OpenStudy (hatahoochee):

Describe how to transform the graph of f into the graph of g. f as a function of x is equal to the square root of x and g as a function of x is equal to the negative square root of x

OpenStudy (hatahoochee):

@kelvin8262

OpenStudy (hatahoochee):

Answers: Reflect the graph of f across the y-axis. Reflect the graph of f across the y-axis and then reflect across the x-axis. Reflect the graph of f across the x-axis. The graph shifts left two units.

OpenStudy (kelvin8262):

so what do you think of f(x) and g(x) what do they equal to?

OpenStudy (hatahoochee):

x is equal to the square root of x

OpenStudy (hatahoochee):

f(x) is equal to the square root of x

OpenStudy (hatahoochee):

g(x)= square root of -x

OpenStudy (hatahoochee):

or -square root of x*****

OpenStudy (kelvin8262):

\[g(x) = -\sqrt{x}\] remember negative value can't be square root. so what does this \[\sqrt{x}\] mean? **look at f(x)

OpenStudy (hatahoochee):

That they're both the same and positive? I'm not really sure

OpenStudy (kelvin8262):

f(x) = -g(x) so imagine you a graph of f(x). now if g(x) is f(x), which mean they are the same. but if there is a negative sign, f(x)=-g(x). what will g(x) look like?

OpenStudy (kelvin8262):